# Intergroup violence in bursts or fizzles

### v9

@@ Abstract
During intergroup confrontations, agitating stimuli such as opponents’ threats and provocations can trigger collective violence, even when the usual mechanisms of cooperation, such as norms with sanctions, are absent. We examine video recordings of street fights between groups of young men. Collective violence in their attacks sometimes breaks out in a burst, while at other times it reaches only a fizzle in which only a few group members participate. An adapted Ising spin-glass model demonstrates that these two temporal unfoldings can be predicted by the proportion of unconditional defectors in a focal group.
@@ 1 Introduction
When observing the unfolding of intergroup violence committed by non-professionals, one notices a burst or a fizzle in the majority of cases. Before punches are thrown, individuals in the focal group and their opponents insult and threaten, and thereby agitate, one another. For this to result in violence by focal group members, they have to be in close proximity and have to have face-to-face contact [15, 14, 36]. When agitation passes a critical level (for example, when one individual is pushed by an opponent), violence can break out in a burst wherein a majority of individuals (or a subgroup from a larger and more dispersed group) participates. Alternatively, the confrontation can fizzle out, even if group members are agitated. In the case of fizzling, one or few individuals push or punch asynchronously, but there is no burst of violence. Additionally, understanding how collective violence unfolds is complicated by other people present on the scene, who may form an audience or try to de-escalate [32, 46, 63]. De-escalation happens often in street fights [47]. Given a common prelude of proximity and agitation mixed with de-escalation, our question is how to explain bursts and fizzles. To this end we use the centennial Ising spin-glass model from physics [55, 44], which resulted in a Nobel Prize for Parisi (2021), and has been used occasionally in the social sciences in various studies of social influence [12, 54, 29, 22]. We apply the model to videos of street fights between groups of (mostly) young men.
@@
Collective violence involves a dilemma of collective action [26, 43, 42], so before predicting bursts and fizzles, we must first address how the Ising model explains overcoming such dilemmas. The face-to-face contacts of the focal group members are the social ties of their network. We already know from public goods experiments that more than half of the participants are conditional cooperators willing to contribute if others do [13], thus conforming to their (weighted) average neighbor in the network [7, 57]. Our empirical study is about relatively inexperienced individuals, rather than seasoned fighters, hence our subjects experience uncertainty when they come under threat, which increases the likelihood of their conforming to the group [65, 39]. Conformity makes sense from an evolutionary perspective when payoffs are hard to predict [59]. Therefore subjects’ intentionality not only depends on their expected payoff (as a share of the public good, in this case attack of or defense against opponents) but also their desire to align with their group members.
@@
Now think of two individuals, each with the behavioral options to defect or cooperate. (Behavioral options correspond to magnetic spins in the original model, but that does not matter here.) The possibilities are: (1) both individuals defect, which avoids exploitation but does not yield any public good; (2) one individual freerides the cooperating other, yielding half of the public good; or (3) both cooperate at a cost, which maximizes the public good. The dashed line in Fig. 1 plots these possibilities, with the normalized number of cooperators ($N_{C}/n$) from left to right. At the beginning, both defect (the local minimum of $H$ on the left). The vertical axis ($H$, or energy in the original) conveys that moving uphill (i.e., that one of the two will start contributing to the public good) is unlikely. By contrast, if moving to lower levels of $H$ is possible, it will certainly happen; thus, if one of the two cooperates, the other conditional cooperator will join in, and the curve goes to the global minimum of $H$, on the right. The hill is the graphical representation of the dilemma of collective action; it is also drawn for a group of five. It shows that without a driving force, it is very unlikely that a group of non-fighting individuals ($n\geq 2$) will overcome the hill and fight collectively, and it is even less likely in larger groups.
@@
Figure 1: The dilemma of collective action presented as a hill between full defection (left) and full cooperation (right), with the proportion of cooperators $(N_{C}/n)$ on the horizontal axis. The vertical axis $(H/n)$ could be intuited as a negative likelihood; moving uphill from a state of defection $(N_{C}/n=0)$ is very unlikely. One line (dashed) is drawn for a dyad and one for a clique of five individuals. The larger the group is, the more rounded the hill becomes.
@@
Nonetheless, some individuals may start fighting, which in turn may prompt others to join. When agitating stimuli, such as opponents’ threats, increase, one or few individuals may accidentally cooperate. This phenomenon of accidental, or spontaneous, cooperation—called “trembling hands” in game theory [17]—can influence proximate others and may result in a cascade of cooperation. For the model, we define turmoil ($T$, or temperature in the original) as opponents’ behavior that agitates members of the focal group. If turmoil increases beyond a critical level ($T_{c}$), fighting can break out in a burst. [^1] The interesting question of why some people self-select into groups where they can fight against others is beyond the scope of this study. Our contribution is to show that the outbreak of collective violence depends on a second critical level, namely the critical proportion of unconditional defectors in the focal group ($p_{c}$). In violent situations, people may defect for several reasons. They may be too scared to fight [15], have empathy with their opponents, disagree with violence (i.e., appraise the public good differently), try to de-escalate, be stopped in their tracks by de-escalators (or guardians), or they may have fought but were wounded, were wrestled to the ground, or became exhausted at some point and became passive. If, for whatever reason, their proportion stays below the critical level ($p\leq p_{c}$), turmoil ($T\geq T_{c}$) will be followed by a burst of violence, but above the critical level ($p>p_{c}$), there is neither a critical level of turmoil nor a burst, and violence occurs in a fizzle. The critical level of steady defectors does not follow from earlier theories, but the Ising model allows us to infer it, which we use as a prediction that we test empirically. Note that $p$ incorporates not only de-escalators in the focal group, but also focal group members who have been prevented from fighting by de-escalators in the focal group or by others.
@@
Most models in the literature revolve around individual rewards and punishments, called selective incentives [42], on top of a share of the public good. These incentives require norms about (in)appropriate behavior in a given situation [19], as well as monitoring of group members [48], and transmission of information (i.e., gossip) through the group’s network, which leads to reputations [41] that feed back through selective incentives, with or without leaders. This package of mechanisms is crucial for ongoing cooperation in the longer run but needs time to develop, which may not be available when threats are imminent. The more time people have, the better they can prepare themselves, which is especially important for high-risk situations. Examples of well-prepared groups are police, soldiers, firefighters, and combat medics who receive professional training that enables them to cooperate effectively and respond to situational stimuli in predetermined manners rather than spontaneously. By contrast, ordinary citizens who face violent opponents are far less prepared, or not prepared at all. For unprepared groups, the uncertainties (of outcomes, benefits, and costs) are higher.
@@
Some earlier models of cooperation do not feature the usual package of cooperation mechanisms; namely, models of thresholds [25], cascades [61], and critical mass [35]. They draw on people who take the initiative or leaders to set cooperation in motion. They also draw on strong rationality assumptions, such as perfect information on the numbers of cooperators. It has been found that in protests, however, leaders often proscribe violence, and most violence occurs in the absence of leaders [27]. Hence, the presence of leaders rarely explains the outbreak of violence. The Ising model, in contrast, points out how cooperation can start spontaneously without leaders, galvanized by accidental cooperators rather than exceptionally zealous ones. [^2] This was already known for the prisoners’ dilemma [17]. If there are initiative takers or leaders [23], however, they can be accommodated, and people may also enter and leave a focal group. The Ising model is also more parsimonious than the aforementioned models because it does not rely on assumptions of strong rationality; the degree of rationality is bounded in violent and other uncertain situations due to incomplete information about opponents and group members beyond face-to-face contact.
@@ 2 Model
Members of a (fledgling) focal group, indexed $i$ or $j$, can defect, $D$, or contribute, $C$, to a public good, with $0<D<C$. Behavioral variable $S_{i}$ of conditional cooperators can take the value $S_{i}=C$ or $S_{i}=-D$, whereas unconditional defectors stay put at $S_{j}=-D$. Before a collective action, everyone in a focal group defects. Network ties among focal group members, $A_{ij}>0$, mean that $i$ is in close proximity to, and senses the behavior of, group member $j$, or else $A_{ij}=0$. Because people tend to respond to proportions of their social environment rather than absolute numbers [61, 21], ties are row-normalized, with $w_{ij}=A_{ij}/\sum_{j=1}^{n}A_{ij}$ such that $\sum_{j=1}^{n}w_{ij}=1$. We assume that sensing is reciprocal to at least some degree but not necessarily symmetrical, and that in our small groups ($n<10$), everyone senses all others, thus the network is fully connected. [^3] Temporary exceptions to full connectedness were groups where some people’s view was blocked by objects, opponents, or bystanders, as well as one larger group ($n=14$). These cases we investigated through simulations. Opponents’ turmoil, represented by $T$, has an effect on the focal group at the aggregate level and affects everyone equally. Locally varying turmoil is discussed elsewhere [9]. The Ising model is the following Hamiltonian equation [5, 60]
@@
We do not assume that individuals know their payoffs in advance, but they will heuristically—and perhaps wrongly—distinguish between valuable ($C>D$) and nonvaluable ($C<D$) public goods or, to the same effect, between potentially efficacious and non-efficacious actions [49]. Note that payoffs are not used in the model’s calculations but are defined to provide a meaningful interpretation. When $i$ chooses between $C$ and $D$ amid $N_{C}$ cooperators, payoffs for cooperation and defection are, respectively
@@
For our empirical study, we have to choose values for $C$ and $D$ to predict the critical level of unconditional defectors, $p_{c}$. The most obvious choice is $C=1$, as in game theory and the original Ising model. For $D$ we want to avoid two trivial values: if $D=1$, there is no point in cooperating, and if $D=0$, there is no dilemma (but a downward slope to the right in Fig. 1). Choosing $D$ to be maximally distant from the trivial values seems to be a reasonable first approximation, hence we set $S=\{1,-1/2\}$ for all conditional cooperators, also in the examples. Because these two values were published earlier [10], one could say they were preregistered. For unconditional defectors we set $S_{j}=-1/2$, irrespective of their reasons. Note that all earlier Ising models had either the values $\{1,-1\}$ [12, 54, 22, 64] or $\{1,0\}$ [16]. Of all these models, only one represents a public goods game, in this case for two individuals [1, 50] whereas our model is applicable to groups of any size.
@@
Solving the Ising model boils down to minimizing $H$, which can be done analytically by a mean field approach (in the Supplementary Material). It abstracts away from the network structure, which is no problem for our fully connected networks, and focuses on the mean behavior. We prove that, if $C=1$ and $D=1/2$, $p_{c}=1/3$. For inhomogeneous (such as clustered) networks, the mean field approach is inaccurate. $H$ can also be minimized computationally, which is much simpler, and can take inhomogeneous networks into account. Here, we explain the computational approach (Fig. 2). For a given level of turmoil, a network node $i$ is randomly picked and $H$ is calculated. For comparison, $i$’s current behavior is flipped, from $D$ to $C$ or the other way around, and $H^{\prime}$ with the flip is calculated. The flip is accepted and implemented if $H^{\prime}<H$ or with a certain chance that increases with $T$. In other words, increasing $T$ increases the amount of randomness (due to situational turmoil) in $i$’s decision. A behavioral change of $i$ affects others in $i$’s neighborhood when it is their turn to decide. Consecutive decisions are Monte Carlo steps in the Metropolis algorithm [5] that loops through great many Monte Carlo steps (counted by $t$ in Fig. 2) in order to allow individuals’ interdependent behavior to settle down. This procedure is repeated at increments of $T$, in our study with step size 0.01.
@@
Figure 2: The Metropolis algorithm. Index $i$ runs over the conditional cooperators (not the unconditional defectors); $t$ is a counter of Monte Carlo steps; $c_{r}$ is a random number in the range $0\leq c_{r}\leq 1$.
Figure 3: Color online. Proportion of cooperators ($N_{C}/n$) with turmoil ($T$) in a fully connected network in the size range of our empirical study ($n=5$); again, $C=1$ and $D=1/2$. The black line (top) depicts the group without unconditional defectors ($p=0$), the red line (middle) a group with one unconditional defector ($p<p_{c}$), and the dashed blue line (bottom) a group with two unconditional defectors ($p>p_{c}$).
@@
To illustrate the emergence of cooperation with increasing turmoil in a group of $n=5$, we apply the Metropolis algorithm Fig. 3. First, there are only conditional cooperators. At low turmoil, collective action does not start (black line at the top), but at a critical level $T_{c}$, (almost) everybody bursts into cooperation, with a maximum (where $N_{C}/n\approx 1$) at or near $T_{c}$. Then, a few accidental cooperators win over most others to join the collective action. The effect of turmoil is nonmonotonic and the level of cooperation decreases if $T$ keeps increasing beyond $T_{c}$, which means that very strong turmoil becomes more confusing than agitating. Fighting ends when exhaustion sets in, a winner stands out (opponents have been wrestled to the ground or have fled), or others intervene.
@@
If there are unconditional defectors, $T_{c}$ increases, which is hardly visible in Fig. 3 but more pronounced in larger networks, and maximum cooperation is lower than in the previous case because the number of conditional cooperators is lower. [^6] In Fig. 1, a proportion of unconditional defectors, $p$, means that no matter how many conditional cooperators contribute, $N_{C}/n\leq 1-p$. If the proportion of unconditional defectors surpasses a critical level, $p_{c}$, there is a fizzle (dashed blue line in Fig. 3) with much lower levels of cooperation, and no burst-like start of cooperation. In the mean field analysis, there is no (burst of) cooperation if $p_{c}>S_{0}/\Delta$, independent of network size and density. In simulations, $p_{c}$ is proportionally less precise in smaller networks (Fig. 4; Table S3), and the prediction for the triad (red dot in Fig. 4) is empirically false whereas the mean field is correct: one unconditional defector does not prevent cooperation of the other two. Also one empirical group of six with two unconditional defectors had a burst, which is consistent with the mean field but did not happen in the simulations. Based on the mean field analysis ($p_{c}=1/3$) and a nearly identical result from the simulations ($p_{c}=0.34$), [^7] If unconditional defectors are clustered together in a sparse network, they are less in the way of collective action of the remainder network (thus $p_{c}$ is higher) than if they are evenly spread out across the network. our main prediction is that $p_{c}=1/3$.
Figure 4: For simulated networks with sizes $3\leq n\leq 100$, the minimum number of unconditional defectors who prevent a burst is plotted. The red dot (left, at the bottom) marks the triad (1 defector prevents a burst of 2), which conflicts with the empirical data (1 defector and a burst of 2).
@@
The critical threshold of turmoil increases with network size at a decreasing rate [9], but it also increases with the proportion of unconditional defectors. At $T_{c}$, simulations point out that cooperation starts in small clusters of conditional cooperators, which we will also test empirically. [^8] This bottom up mounting of cooperation is similar to bottom up synchronization [3]. This outcome is puzzling in real life because larger groups have a better chance to win at lower individual costs. Yet if in a small (sub)group, someone starts fighting, he accounts for a relatively large proportion of his neighbors’ social contacts, and more readily wins them over to fight than in a large group where he would comprise a small minority.

### v10

@@ Abstract
During intergroup confrontations, agitating stimuli such as opponents’ threats and provocations can trigger collective violence, even when the usual mechanisms of ingroup cooperation, such as norms with sanctions, are absent. We examine video recordings of street fights between groups of young men. Their violence sometimes breaks out in a burst, while at other times it reaches only a fizzle in which only a few group members participate. An adapted Ising model explains the dilemma of collective action and demonstrates that these two temporal unfoldings can be predicted by the proportion of defectors in a focal group.
@@ 1 Introduction
In studies of intergroup conflict, the large numbers of casualties of wars between and within countries draw most attention [17, 59]. Yet, the most frequent occurrences of intergroup violence involve small groups, including subgroups of larger groups. We focus on small groups, and ask how violence committed by non-professionals unfolds. When observing pertinent videos, one notices a burst or a fizzle in the majority of cases. Before punches are thrown, individuals in the focal group and their opponents insult and threaten, and thereby agitate, one another. For this to result in violence by focal group members, they have to be in close proximity and have to have face-to-face contact [19, 16, 45].
In one of the most influential studies of violence, Randall Collins [19] argues that antagonists need to circumvent an emotional barrier of tension and fear in order to be able to use violence. Surmounting the barrier requires that assailants attain a state of emotional dominance. In face-to-face interpersonal conflicts, they can reach emotional dominance when the other party appears vulnerable or weaker, or when the focal group is supported by an audience [19]. Collins’ notion of “forward panics” corresponds to the bursts we observed. In forward panics, assailants seize the opportunity to forcefully attack opponents in moments of weakness or vulnerability, for example when being isolated or falling down to the ground. This emotional barrier raises another barrier known as the dilemma of collective action [53, 54], which has gained less attention in scholarly work on violence. We will present a formal model that explicates the conditions for groups to burst into collective violent action. At the core of the model is opponents’ agitation.
@@
When agitation passes a critical level (for example, when one individual is pushed by an opponent), violence can break out in a burst wherein a majority of individuals (or a subgroup from a larger and more dispersed group) participates. Alternatively, the confrontation can fizzle out, even if group members are agitated. In the case of fizzling, one or few individuals push or punch asynchronously, but there is no burst of violence. Additionally, understanding how collective violence unfolds is complicated by other people present on the scene, who may form an audience or try to de-escalate [39, 57, 79]. De-escalation happens often in street fights [58]. Given a common prelude of proximity and agitation mixed with de-escalation, our question is how to explain bursts and fizzles. To this end we use the centennial Ising model from physics [43, 70]. It resulted in a Nobel Prize for Parisi (in 2021; [55]), and has been used in various studies of social influence [14, 69, 34, 26]. We adapt the model in a novel way, use it as an agent based model, and apply it to videos of street fights between groups of (mostly) young men.
2 Theory
@@
Collective violence involves a dilemma of collective action [54, 53], so before predicting bursts and fizzles, we must first address how the Ising model explains overcoming such dilemmas. The face-to-face contacts of the focal group members are the social ties of their network. We already know from public goods experiments that if participants regard a public good (in our case the attack of or defense against opponents) to be valuable, more than half of the participants are conditional cooperators willing to contribute if others do [15]. In other words, they conform to their (weighted) average neighbor in the network [7, 73]. Our empirical study is about relatively inexperienced individuals, rather than seasoned fighters, hence our subjects experience uncertainty when they come under threat, which increases the likelihood of their conforming to the group [81, 49]. Conformity makes sense from an evolutionary perspective when payoffs are hard to predict [75]. Therefore, subjects’ motivation not only depends on their appraisal of the public good, but also their desire to align with their group members. [^1] The interesting question of why certain people self-select into groups where they can fight against others is beyond the scope of this study. Milliff [47] suggested that these people perceive their situation as uncertain but controllable, which is consistent with our assumptions.
@@
Now think of two individuals, each with the behavioral options to defect or cooperate. Behavioral options correspond to magnetic spins in the original Ising model, but that is not relevant here; we apply the model differently and do not establish an analogy with the original. The possibilities are: (1) both individuals defect, which avoids exploitation but does not yield any public good; (2) one individual freerides the cooperating other, yielding half of the public good; or (3) both cooperate at a cost, which maximizes the public good. The dashed line in Fig. 1 plots these possibilities, with the normalized number of cooperators ($N_{C}/n$) from left to right. At the beginning, both defect (the local minimum of $H$ on the left). The vertical axis ($H$, or energy in the original) conveys that moving uphill (i.e., that one of the two will start contributing to the public good) is unlikely. After all, most individuals understand that fighting alone against a group is very dangerous. By contrast, if moving to lower levels of $H$ is possible, it will certainly happen; thus, if one of the two cooperates, the other conditional cooperator will join in, shifting $H$ to its global minimum on the right. The hill is the graphical representation of the dilemma of collective action; it is also drawn for a group of five. It shows that without a motivating force, it is very unlikely that a group of non-fighting individuals ($n\geq 2$) will overcome the hill and fight collectively, and it is even less likely in larger groups.
@@
Figure 1: The dilemma of collective action presented as a hill between full defection (left) and full cooperation (right), with the proportion of cooperators $(N_{C}/n)$ on the horizontal axis. Data points are based on Equation 1, for $C$ = 1 and $D$ = 1/2. The vertical axis $(H/n)$ could be intuited as a negative likelihood; moving uphill from a state of defection $(N_{C}/n=0)$ is very unlikely. One line (dashed) is drawn for a dyad and one for a clique of five individuals. The larger the group is, the more rounded the hill becomes.
Nonetheless, some individuals may start fighting, which in turn may prompt others to join. When agitating stimuli, such as opponents’ threats, increase, focal group members become aroused and angry. Consequently, one or few individuals may accidentally cooperate even when most others do not (yet). This phenomenon of accidental, or spontaneous, cooperation—called “trembling hands” in game theory [21]—can influence proximate others and may result in a cascade of cooperation.
@@
For the model, we define turmoil ($T$, or temperature in the original) as opponents’ behavior that agitates members of the focal group, which (in our case) means that their arousal and anger increase; the hormones adrenaline and cortisol are produced and linger on for a while. In the model, opponents’ turmoil has an effect on the focal group at the aggregate level and affects everyone equally. In larger groups than ours, individuals standing closer to opponents are exposed to more turmoil than others standing far away. Locally varying turmoil in the Ising model is discussed elsewhere [10]; here, we avoid unnecessary complications. When keeping it simple, the model demonstrates that if turmoil increases, there is no gradual increase of accidental cooperators. Instead, when turmoil reaches a critical level ($T_{c}$), fighting breaks out in a burst; see Figure 2. Our contribution is to show that the outbreak of collective violence depends on a second critical level, namely the critical proportion of unconditional defectors in the focal group ($p_{c}$). In violent situations, many people do not fight, for all kinds of reasons. They may be too scared to fight [19], have empathy with their opponents, disagree with violence (i.e., appraise the public good differently), feel no solidarity with their group [18], try to de-escalate, be stopped in their tracks by de-escalators (also called guardians), or they may have fought but were wounded, were wrestled to the ground, or became exhausted at some point and became passive. If their proportion stays below the critical level ($p\leq p_{c}$), turmoil ($T>T_{c}$) will be followed by a burst of violence. Above the critical level ($p>p_{c}$), however, there is neither a critical level of turmoil nor a burst, and violence occurs in a fizzle (Fig. 2). The critical level of unconditional defectors does not follow from earlier theories, but the Ising model allows us to infer it, which we use as a prediction that we test empirically. Note that $p$ incorporates not only de-escalators in the focal group, but also focal group members who have been prevented from fighting by de-escalators in the focal group or by others.
Figure 2: Color online. Simulation of Equation 1 over a range of turmoil ($T$) on the horizontal axis. The proportion of cooperators ($N_{C}/n$) in a fully connected network in the size range of our empirical study ($n=5$) is on the vertical axis; $C=1$ and $D=1/2$. The black line (top) depicts a burst in the group without unconditional defectors ($p=0$), the red line (middle) a burst in the group with one unconditional defector ($p<p_{c}$), and the dashed blue line (bottom) a fizzle in the group with two unconditional defectors ($p>p_{c}$).
@@
Most models of collective action in the literature revolve around individual rewards and punishments, called selective incentives [53], on top of a share of the public good. These incentives require norms about (in)appropriate behavior in a given situation [23], as well as monitoring of group members [61], and transmission of information (i.e., gossip) through the group’s network, which leads to reputations [51] that feedback through selective incentives, with or without leaders. This package of mechanisms is crucial for ongoing cooperation in the longer run but needs time to develop, which may not be available when threats are imminent. The more time people have, the better they can prepare themselves, which is especially important for high-risk situations. Examples of well-prepared groups are police, soldiers, firefighters, and combat medics who receive professional training that enables them to cooperate effectively and respond to situational stimuli in predetermined manners rather than spontaneously. By contrast, ordinary citizens who face violent opponents are far less prepared, or not prepared at all. For non-professional groups, such as ours, the uncertainties (of outcomes, benefits, and costs) are higher.
@@
Several earlier models of cooperation do not feature the usual package of cooperation mechanisms; namely, models of thresholds [30], cascades [77], and critical mass [44]. Therefore, these models are potentially useful for our study. They draw on the assumption that some people take the initiative or leaders set cooperation in motion. It has been found that in protests, however, leaders often proscribe violence, and most violence occurs in the absence of leaders [32]. Hence, the presence of leaders does not explain the outbreak of violence in general. The Ising model, in contrast, points out how cooperation can start spontaneously without leaders, galvanized by accidental cooperators rather than exceptionally zealous ones. If there are initiative takers or leaders in favor of violence [28], however, they can be accommodated in our model, and people may also enter and leave a focal group.
These models also draw on strong rationality assumptions, such as perfect information on the numbers of cooperators at every moment. Yet, the degree of rationality is bounded in violent and other uncertain situations due to incomplete information about opponents and group members beyond face-to-face contact. Moreover, inexperienced fighters cannot assess their payoffs with any degree of accuracy before a fight. The Ising model is more parsimonious than the aforementioned models because it does not rely on assumptions of strong rationality. Individuals may have fairly accurate expectations of their payoffs, but for the model it is enough if they find the public good valuable or not. The model combines the precision of game theory, including a mapping on payoffs—unknown to the participants—with the behavioral spontaneity of chaotic and uncertain situations [71], which is usually assessed qualitatively [68], as well as a portion of randomness in individuals’ decisions, reflecting their uncertainty [42]. Moreover, whereas game theoretic approaches to conflict readily become complicated, even for two individuals [64], the simplicity of the Ising model makes it possible to investigate large networks computationally.
Sociology’s most influential, qualitative model of violence has been developed by Collins [19]. Our work contributes to his model. One of its important contributions was that it shifted researchers’ attention to what actually happens in fights, for which they started using visual data (e.g., [50]). At the core, his model has potentially quantifiable variables capturing emotions, such as confrontational tension (agitation in our model), forward panic (our burst), and emotional dominance, to predict the beginning of violence and the winner in the end. Empirical studies using these variables have remained qualitative [8, 31, 36], to which we add different, quantitative insights. We build on Collins’ work, also use visual data, but our general strategy is using behavioral variables that are relatively easily observable. We minimize the number of variables from the literature that are hard to observe behaviorally, such as reputations, norms, emotions, and solidarity, without downplaying their importance; they can be incorporated in an extended version of the Ising model [9] but we lack video recorded data to include them in our study. Complementing Collins, our approach is explicitly relational, adds precision, and explains collective violent action. We will first demonstrate how the Ising model explains surmounting collective action’s barrier.
@@ 3 Model
Members of a (fledgling) focal group, indexed $i$ or $j$, can defect, $D$, or contribute, $C$, to a public good, with $0<D<C$. Behavioral variable $S_{i}$ of conditional cooperators can take the value $S_{i}=C$ or $S_{i}=-D$, whereas unconditional defectors stay put at $S_{j}=-D$. Before a collective action, everyone in a focal group defects. Network ties among focal group members, $A_{ij}>0$, mean that $i$ is in close proximity to, and senses the behavior of, group member $j$, or else $A_{ij}=0$. Because people tend to respond to proportions of their social environment rather than absolute numbers [77, 25], ties are row-normalized, with $w_{ij}=A_{ij}/\sum_{j=1}^{n}A_{ij}$ such that $\sum_{j=1}^{n}w_{ij}=1$. Based on our video data, we assume that in out small groups ($n<10$), sensing is reciprocal at least to some degree, but not necessarily symmetrical, and that everyone senses all others, thus the network is fully connected. [^2] Temporary exceptions to full connectedness in our data were groups where some people’s view was blocked by objects, opponents, or bystanders, as well as one larger group ($n=14$). These cases we investigated through simulations. The model can be used for complex networks as well, beyond the small groups studied here [10].
The Ising model is the following Hamiltonian equation [5, 76]
@@
We do not assume that individuals know their payoffs in advance, but they will heuristically—and perhaps wrongly—distinguish between valuable ($C>D$) and unvaluable ($C<D$) public goods or, to the same effect, between potentially efficacious and non-efficacious actions [62]. Note that payoffs are not used in the model’s calculations but are defined to provide a meaningful interpretation. When $i$ chooses between $C$ and $D$ amid $N_{C}$ cooperators, payoffs for cooperation and defection are, respectively
@@
For our empirical study, we have to choose values for $C$ and $D$ to predict the critical level of unconditional defectors, $p_{c}$. The most obvious choice is $C=1$, as in game theory and the original Ising model. For $D$ we want to avoid two trivial values: if $D=1$, there is no point in cooperating, which is equally valuable as defecting, and if $D=0$, there is no dilemma (but a downward slope to the right in Fig. 1) and violence becomes self-explanatory. Choosing $D$ to be maximally distant from the two trivial values seems to be a reasonable first approximation. Hence, we set $S=\{1,-1/2\}$ for all conditional cooperators, also in the examples. These two values were published in a theoretical study preceding this empirical study [11]. For unconditional defectors we set $S_{j}=-1/2$, irrespective of their reasons, simply because they all discourage the conditional cooperators to fight (Fig. 1). Unconditional means under the conditions in our cases; the notion juxtaposes them to conditional cooperators. Note that all earlier Ising models had either the values $\{1,-1\}$ [14, 69, 26, 80] or $\{1,0\}$ [20]. Of all these models, only one represents a public goods game, in this case for two individuals [1, 63] whereas our model is applicable to groups of any size.
@@
Solving the Ising model boils down to minimizing $H$, which can be done analytically by a mean field approach (in the Supplementary information). We prove that, if $C=1$ and $D=1/2$, $p_{c}=1/3$. For inhomogeneous (such as clustered) networks, the mean field approach is inaccurate, but $H$ can also be minimized computationally, which is simpler and can deal with inhomogeneous networks.
@@
The computational approach, over a certain range of $T$, is as follows (Fig. 3). For a given level of turmoil, a network node $i$ is randomly picked and $H$ is calculated. For comparison, $i$’s current behavior is flipped, from $D$ to $C$ (or the other way around if $i$ cooperates), and $H^{\prime}$ with the flip is calculated. The flip is accepted and implemented if $H^{\prime}<H$ or with a certain chance that increases with $T$. In other words, increasing $T$ increases the amount of randomness (due to situational turmoil) in $i$’s decision. A behavioral change of $i$ affects others in $i$’s neighborhood when it is their turn to decide. Consecutive decisions are Monte Carlo steps in the Metropolis algorithm [5] that loops through great many Monte Carlo steps (counted by $t$ in Fig. 3) in order to allow individuals’ interdependent behavior to settle down. This procedure is repeated at increments of $T$, in our study with step size 0.01.
@@
Figure 3: The Metropolis algorithm. Index $i$ runs over the conditional cooperators, not the unconditional defectors; $t$ is a counter of Monte Carlo steps; $c_{r}$ is a random number in the range $0\leq c_{r}\leq 1$.
@@
To illustrate the emergence of cooperation with increasing turmoil in a group of $n=5$, we apply the Metropolis algorithm; see Figure 2. First, there are only conditional cooperators. At low turmoil, collective action does not start (black line at the top), but at a critical level $T_{c}$, (almost) everybody bursts into cooperation, with a maximum (where $N_{C}/n\approx 1$) at or near $T_{c}$. Then, a small number of accidental cooperators win over most others to join the collective action. The effect of turmoil is nonmonotonic and the level of cooperation decreases if $T$ keeps increasing beyond $T_{c}$, which means that extraordinarily strong turmoil becomes more confusing than motivating to fight.
@@
If there are unconditional defectors, $T_{c}$ increases, which is more pronounced in larger networks than in Figure 2, and maximum cooperation is lower than in the previous case because the number of conditional cooperators is lower. [^5] In Fig. 1, a proportion of unconditional defectors, $p$, means that no matter how many conditional cooperators contribute, $N_{C}/n\leq 1-p$. If the proportion of unconditional defectors surpasses a critical level, $p_{c}$, and $T$ increases above $T_{c}$, there is a fizzle (dashed blue line in Fig. 2) with much lower levels of cooperation, and no burst-like start of cooperation. In the mean field analysis, there is no (burst of) cooperation if $p_{c}>S_{0}/\Delta$, independent of network size and density. In simulations, $p_{c}$ is proportionally less precise in smaller networks (Table S3), but because our networks are complete, we stick to the mean field prediction. Based on the mean field analysis ($p_{c}=1/3$) and a nearly identical result from the simulations ($p_{c}=0.34$), [^6] If unconditional defectors are clustered together in a sparse network, they are less in the way of collective action of the remainder network (thus $p_{c}$ is higher) than if they are evenly spread out across the network. our main prediction is that $p_{c}=1/3$.
@@
The critical threshold of turmoil increases with network size at a decreasing rate [10], but it also increases with the proportion of unconditional defectors. At $T_{c}$, simulations point out that cooperation starts in small clusters of conditional cooperators, which we will also test empirically. The onset in small groups is puzzling because larger groups have a better chance to win at lower individual costs. Yet if in a small (sub)group, someone starts fighting, he (and rarely she) accounts for a relatively sizeable proportion of his neighbors’ social contacts, and more readily wins them over to fight than in a large group where he would comprise a small minority.
Fighting ends when exhaustion sets in, a winner stands out (opponents have fled or have been wrestled to the ground), or others intervene.